Find the length of the sides of a right triangle if one leg is 1cm longer than the other leg, and the hypotenuse is 2cm longer than the shorter leg.
Answer by richwmiller(17219)

a^2+(a+1)^2=(a+2)^2
My guess without solving is 3,4, 5 right triangle.
a^2+a^2+2a+1=a^2+4a+4
a^2=2a+3
a^2-2a-3=0
(a-3)*(a+1)=0
a=3
a+1=4
a+2=5
So it is the 3,4,5 right triangle.